# DanceChat: Large Language Model-Guided Music-to-Dance Generation

### v1

@@ Abstract.
Music-to-dance generation aims to synthesize human dance motion conditioned on musical input. Despite recent progress, significant challenges remain due to the semantic gap between music and dance motion, as music offers only abstract cues, such as melody, groove, and emotion, without explicitly specifying the physical movements. Moreover, a single piece of music can produce multiple plausible dance interpretations. This one-to-many mapping demands additional guidance, as music alone provides limited information for generating diverse dance movements. The challenge is further amplified by the scarcity of paired music and dance data, which restricts the model’s ability to learn diverse dance patterns. In this paper, we introduce DanceChat, a Large Language Model (LLM)-guided music-to-dance generation approach. We use an LLM as a choreographer that provides textual motion instructions, offering explicit, high-level guidance for dance generation. This approach goes beyond implicit learning from music alone, enabling the model to generate dance that is both more diverse and better aligned with musical styles. Our approach consists of three components: (1) an LLM-based pseudo instruction generation module that produces textual dance guidance based on music style and structure, (2) a multi-modal feature extraction and fusion module that integrates music, rhythm, and textual guidance into a shared representation, and (3) a diffusion-based motion synthesis module together with a multi-modal alignment loss, which ensures that the generated dance is aligned with both musical and textual cues. Extensive experiments on AIST++ and human evaluations show that DanceChat outperforms state-of-the-art methods both qualitatively and quantitatively. The additional qualitative samples can be found at https://dancechat.github.io/anon/.
3D Motion Synthesis, Cross-Modality Learning, Tempo Synchronization
ccs: Computing methodologies Motion processingccs: Applied computing Sound and music computing
Figure 1. Our approach inspired by real-world dance learning. Just as dance learners rely on choreographers to interpret music, DanceChat uses textual dance instructions to bridge the semantic gap between music and motion. Large Language Models (LLMs) act as pseudo-choreographers, transforming musical cues into detailed, actionable guidance for dance generation.
1. Introduction
@@
Dance has long served as a powerful medium of expression in cultures, playing a central role in rituals, social connections, and artistic activities (Carroll and Moore, 2008). In the digital age, dance flourishes on social media, allowing people to share creativity and connect globally. Recent advances in deep learning have enabled computational exploration of this art form, particularly through music-to-dance generation, where models automatically create dance movements synchronized to music. This technology has impactful applications in entertainment, enabling realistic, rhythmically aligned virtual idols and 3D characters in virtual reality. Beyond entertainment, automated dance generation is valuable in fields like robotics, digital human modeling, and physical rehabilitation, where it helps create lifelike movement simulations and supports human-computer interaction.
@@
Most existing music-to-dance generation approaches fall into two main categories: regression models (Huang et al., 2021; Li et al., 2021), which are typically designed to directly learn a mapping from input music to dance motions, and generative models (Siyao et al., 2022; Tseng et al., 2023; Zhang et al., 2024b; Huang et al., 2024), which aim to learn a distribution over plausible dances conditioned solely on the input music. However, both approaches lack explicit movement instructions, and rely only on abstract nature of musical elements such as rhythm, tempo, and mood. This creates a broad semantic gap between music and motion, as illustrated in Fig. 1, and complicates the task of generating physically plausible movements that align with the music’s genre and maintain aesthetic quality. Moreover, the scarcity of labeled dance data, due in part to the high cost of motion capture, makes it difficult for models to learn a well-generalized and diverse motion distribution. With limited training samples, the model struggles to capture the full variability of human dance, resulting in outputs that may lack stylistic diversity or fail to generalize to unseen music.
@@
These challenges highlight the need for advanced methods capable of bridging complex semantic gaps and injecting semantic-rich textual instructions without relying on additional annotations. Recent advancements in large language models (LLMs) have demonstrated remarkable success in capturing nuanced semantic relationships across various domains (Wu et al., 2023; Han et al., 2023; Doh et al., 2024; Gardner et al., 2024; Ding et al., 2024). They are trained on diverse content, including detailed discussions on music, enabling them to interpret musical elements such as tempo, key, and chord progressions. This allows them to understand and generate contextually appropriate descriptions that capture the structure, rhythm, and mood of a piece. This inspires new approaches for tasks involving multi-modal relationships and encourages us to investigate the possibilities for leveraging LLMs to enhance the controllability of music-to-dance generation.
@@
Building on these insights, our method adapts the capabilities of LLMs to address the semantic gap in music-driven dance generation. We propose a novel approach, named DanceChat, which conceptualizes the generation model as a dancing learner and leverages LLMs as a pseudo-choreographer. Specifically, we extract structured musical descriptors, including key, tempo, and chord progressions alongside the dance genre context, and feed this information into LLM. The model then generates corresponding textual dance instructions, which serve as a pseudo-guide to condition the motion generation model during training. This helps reduce the semantic gap by offering interpretative cues that map abstract musical elements to concrete motion patterns.
@@
To further enhance alignment between modalities, we encode the input music into finer-grained descriptors such as beat and downbeat positions, which are known to correlate strongly with movement timing in dance. These rhythmic cues, together with the LLM-generated textual instructions, are integrated through a multi-modal fusion module, enabling the model to form a richer and more coherent understanding of the music. By combining musical structure, rhythmic signals, and semantic instructions, DanceChat produces dance motions that are rhythmically aligned, physically plausible, and stylistically expressive.
@@ 2. Related Work
Figure 2. Overview of our approach. DanceChat consists of three main components: (a) Pseudo Instruction Generation module extracts music caption (tempo, key, chord, etc.) from the given music condition, (b) Multi-modal Fusion module integrates the encoded representations of music, beats, and text into a unified conditional embedding, and (c) Diffusion-based Motion Synthesis module uses this unified embedding to guide the generation of realistic dance movement.
2.1. 3D Human Motion Generation
Common approaches to 3D human motion generation (Zhu et al., 2023) can be broadly categorized into three sub-tasks: text-conditioned motion generation (Ahn et al., 2018; Tevet et al., 2022a; Chen et al., 2023; Zhang et al., 2022; Kim et al., 2023; Lin et al., 2023; Zhang et al., 2023; Zhou and Wang, 2023), which generates motions corresponding to natural language descriptions; scene-conditioned motion generation (Araújo et al., 2023; Ghosh et al., 2023; Huang et al., 2023; Wang et al., 2022), which produces plausible human motions in relation to the surrounding scene layout; and audio-conditioned motion generation, which includes speech-to-gesture generation for conversational scenarios (Yang et al., 2023a; Alexanderson et al., 2023; Ao et al., 2023) as well as music-to-dance generation, which will be discussed in detail later.
Among these, text-conditioned motion generation has received growing attention due to the expressive power and clarity of natural language. In particular, diffusion-based approaches have recently achieved state-of-the-art performance in this area. For example, MotionDiffuse(Zhang et al., 2022) demonstrates that free-form textual input can guide the generation of diverse, coherent, and stylistically rich motion sequences. Extensions such as FLAME(Kim et al., 2023), BEing-Diffusion(Lin et al., 2023), and T2M-GEN(Zhang et al., 2023) incorporate fine-grained control over motion categories, temporal structure, and global style. UDE (Zhou and Wang, 2023) further introduces a unified representation to support cross-task generalization and improved semantic grounding. These methods demonstrate the effectiveness of using natural language as a precise, interpretable, and semantically expressive modality for guiding human motion generation.
2.2. Music-to-Dance Generation
The Music-to-Dance generation task poses unique challenges due to the need for models to understand and generate dance sequences that are both rhythmically and stylistically aligned with music. Previous methods (Chan et al., 2019; Zhuang et al., 2022; Li et al., 2023; Tseng et al., 2023; Yang et al., 2023b; Zhang et al., 2024b; Huang et al., 2024) incorporate advanced audio-processing techniques with motion generation algorithms to enhance the flexibility and expressiveness of generated dances. EDGE (Tseng et al., 2023) employs Jukebox for enriched music audio representation and introduces a contact consistency loss to reduce physical implausibilities.
@@ 2.2. Music-to-Dance Generation
Despite these advancements, existing approaches still face limitations in bridging the substantial semantic gap between music and dance, making the synchronization and aesthetic alignment of dance sequences a significant challenge. Unlike text-to-motion generation (Zhang et al., 2024a; Zhong et al., 2023), where textual cues explicitly describe body movements, music provides only implicit cues through rhythm, tempo, and tone, which do not directly correspond to physical actions. This gap complicates the creation of dance sequences that are both well-aligned with the music and stylistically diverse.
2.3. Large Language Models
Recent instruction-tuned LLMs, such as ChatGPT (OpenAI, 2023), Claude (Anthropic, 2023), LLaMA (Touvron et al., 2023), and DeepSeek (DeepSeek, 2024) have demonstrated strong capabilities in capturing contextual semantics, following user intent, and generalizing across domains. Inspired by these advancements, researchers have begun exploring LLMs for non-linguistic tasks, including motion generation (Zhou et al., 2024; Ribeiro-Gomes et al., 2024), where their ability to model temporal coherence and semantic structure enables fine-grained motion control.
@@ 2.3. Large Language Models
To effectively integrate LLMs into non-linguistic domains, prompt engineering has emerged as a key technique (Liu et al., 2021; Liu and Chilton, 2022; Oppenlaender, 2023). This involves crafting structured prompts that effectively communicate with models, directing the generation toward desired outcomes, and allowing users to control both the content and style of the outputs. In motion generation, prompt engineering enables models to translate high-level textual instructions into semantically coherent motion sequences, enhancing controllability and expressiveness (Wang et al., 2024). Building upon these developments, our work introduces a novel approach that leverages LLM-generated textual choreography prompts to bridge the semantic gap between music and dance.
3. Our Approach
3.1. Problem Definition
@@
The task of music-to-dance generation involves producing a dance motion sequence $\mathcal{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{N}\}$ from a given music track $\mathcal{M}$ with a duration of $\mathit{N}$ frames. Here, $\mathbf{x}_{i}\in\mathbb{R}^{D}$ represents a human pose at the i-th frame, which is denoted as a D-dimensional vector. Following the pose representation in EDGE (Tseng et al., 2023), we present the dance sequences with 24-joint SMPL format (Loper et al., 2023), which represents each joint with 6-DOF rotation representation (Zhou et al., 2019) and a single root translation. Along with the binary contact labels $b\in\{0,1\}^{2\times 2=4}$ for the heel and the toe of each foot, the dimension $D=24\times 6+3+4=151$.
3.2. Overview
@@
Our model architecture is illustrated in Fig. 2 and consists of three main components: Pseudo Instruction Generation, Multi-modal Feature Extraction and Fusion, and Diffusion-based Motion Synthesis. The Pseudo Instruction Generation module (Fig. 2a) extracts music caption (tempo, beat, key, chord) from the given music condition $\mathcal{M}$ and utilizes an LLM as a pseudo-choreographer to transform this information into dance motion instructions $\mathcal{T}$, providing initial guidance for motion generation. The Multi-modal Fusion module (Fig. 2b) then integrates the encoded representations of music $\mathcal{M}$, beats $\mathcal{B}$, and text $\mathcal{T}$ into a unified conditional embedding. This fused representation aligns music, beat, and movement intentions, creating a cohesive context for dance generation. Finally, the Diffusion-based Motion Synthesis component (Fig. 2c) uses this unified embedding to guide the generation of realistic, rhythmically aligned dance movements that capture the style and dynamics of the music, producing smooth and expressive dance sequences.
3.3. Pseudo Instruction Generation
Figure 3. Details of our textual instruction generation process. We construct prompts using a templated format. The prompt guides an LLM to generate HumanML3D-style dance instructions with constraints on motion focus, perspective, and length. The bottom section shows an example of a prompt and its corresponding LLM-generated instruction.
@@ 3.3. Pseudo Instruction Generation
This module is responsible for transforming music into high-level dance instructions, which serve as semantic guidance for the dance generation process. This module comprises two key steps: Music Caption Construction and Textual Instruction Generation.
Music Caption Construction. We extract essential musical attributes from the input music $\mathcal{M}$ to create a comprehensive music caption. Following Mustango (Melechovsky et al., 2023), we utilize multiple specialized tools to carry out the extraction:
  • Tempo: We employ BeatNet (Heydari et al., 2021) to extract tempo information, measured in beats per minute (BPM), to capture the pace of the music. This tempo provides a foundational cue for the timing of generated dance movements.
  • Chord Progression: Using Chordino (Mauch and Dixon, 2010), we identify the sequence of chords within the music. Chord progression conveys harmonic changes, which influence the emotional tone of the dance.
  • Key: The key of the music is extracted using Essentia’s (Bogdanov et al., 2013) KeyExtractor, which determines the tonal center of the piece, providing coherence and structure to the generated dance.
  • Beat Structure: BeatNet (Heydari et al., 2021) is further utilized to detect beats and downbeats, which serve as rhythmic anchors for synchronizing dance movements with the music.
Textual Instruction Generation. Given the constructed music caption, we generate a textual dance instruction $\mathcal{T}$ that aligns with the corresponding music clip $\mathcal{M}$. The instruction serves as a pseudo-guidance for training and is produced using a structured prompt directed to an LLM, which acts as a pseudo choreographer. To ensure that the generated instruction is rhythmically and stylistically appropriate, we adopt a carefully designed prompt engineering strategy inspired by the format and motion description style of the HumanML3D dataset (Guo et al., 2022), a comprehensive collection comprising diverse text-motion pair data. As shown in Fig 3, we designed a prompt engineering strategy that incorporates several key considerations to guide the model’s output effectively. First, we include the duration of the music clip in the prompt, ensuring that generated instructions are aligned with the length and pacing of the music. Additionally, to produce genre-specific movement, the prompt specifies a dancing style (e.g., hip-pop, ballet) extracted from data labels, providing the model with style supervision to create dance instructions that match both the music’s rhythm and genre. To reduce noise in the generated text, we explicitly direct the model to avoid directional words (e.g., left, right), which can introduce ambiguity or irrelevant details. This focus on general body movements without spatial directions helps maintain clarity and relevance in the instructions. We also aim to control the text style, structuring the prompt concisely and descriptively.
@@
Instruction Results. The output of this process is a set of natural language dance instructions that provide semantic cues for the dance generation model. For instance, given a music track with a moderate tempo and a hip-hop style, the generated instruction might read: The dancer bounces lightly to the beat, alternating shoulder pops with a subtle chest pump. These instructions capture both the rhythmic alignment and stylistic characteristics of the music, bridging the semantic gap between audio and motion. Example is shown in Fig. 3.
3.4. Multi-modal Feature Extraction and Fusion
3.4.1. Multi-modal Feature Extraction
Music Encoding. We leverage a pre-trained Jukebox (Dhariwal et al., 2020) model to extract high-level musical features from raw audio inputs. Specifically, we input the raw audio into the frozen Jukebox encoder, obtaining embeddings that represent the musical content. To adapt these embeddings for our task, we introduce a trainable two-layer Transformer encoder that processes the Jukebox-derived features, allowing the model to learn task-specific representations while retaining the rich musical information captured by Jukebox, resulting in embeddings $E_{M}\in\mathbb{R}^{d_{M}}$.
Text Encoding. The dance instruction generated by LLMs is processed through a CLIP (Radford et al., 2021) text encoder to capture semantic details about body movements and style. To tailor the text embeddings to our specific application, we incorporate a learnable 4-layer Transformer as a conditioning encoder. This additional Transformer processes the frozen CLIP embeddings, enabling the model to capture extra semantic dance guidance. To better leverage prior knowledge from existing text-to-motion models, we initialize our entire text encoder with the pretrained weights from MotionDiffuse (Zhang et al., 2022). This is motivated by the fact that our pseudo instructions are generated to closely follow the style and vocabulary of the HumanML3D (Guo et al., 2022) dataset, which was also used to train MotionDiffuse. Since our textual instructions are automatically generated and no paired text-motion supervision is available, we do not apply direct supervised training to the text encoder. Instead, we rely on the strong prior learned from MotionDiffuse to provide a semantically meaningful representation. The pretrained CLIP and Transformer modules together form our text encoder, and the final textual embedding is denoted as $E_{T}\in\mathbb{R}^{d_{T}}$.
Beat Encoding. In addition to basic musical features, we explicitly incorporate beat proximity information to enhance rhythmic awareness in generated motion. We apply Librosa (McFee et al., 2015) to extract both beat and downbeat positions from the audio. These positions are discrete time points that mark the perceived rhythmic structure of the music. To represent rhythmic proximity in a continuous form, we follow a Gaussian-based approach: for each frame, we compute a Gaussian-like pulse centered around the nearest beat and downbeat positions. This encoding reflects the temporal distance to nearby rhythmic accents, enabling the model to perceive and respond to the rhythmic context over time. We then project these proximity values into a learnable embedding space using a linear projection layer, resulting in beat embeddings $E_{B}\in\mathbb{R}^{d_{B}}$.
Together, these extracted features capture the unique attributes of each modality, providing a foundation for effective cross-modal integration. Each of these extracted feature embeddings provides distinct complementary information: textual embedding $E_{T}$ offers semantic pseudo guidance, music embedding $E_{M}$ captures musical style, and beat embedding $E_{B}$ encodes rhythmic timing.
3.4.2. Multi-modal Hierarchical Fusion
Musical-Rhythmic Fusion. Initially, we fuse music embeddings $E_{M}$ with rhythmic beat embedding $E_{B}$ using an additive fusion approach:
(1) $E_{M}^{\prime}=E_{M}+E_{B},$
Music-Text Semantic Fusion. Subsequently, we incorporate textual information from dance instruction embeddings $E_{T}$ to provide semantic guidance for dance generation. Specifically, we concatenate the textual embeddings $E_{T}$ with the previously fused musical-rhythmic embeddings $E_{M}^{\prime}$:
(2) $E_{F}=[E_{M}^{\prime};E_{T}],$
The final conditional embedding $E_{F}$ encapsulates a balanced representation of the text, music, and beat, ensuring that generated dance movements are not only synchronized with the rhythm but also stylistically aligned and semantically meaningful. This fused embedding serves as the input for the diffusion-based motion synthesis module, enabling the generation of realistic, expressive dance sequences that reflect the musical nuances and choreographic intentions embedded in the text.
3.5. Diffusion-based Motion Synthesis
This module leverages a denoising diffusion probabilistic model (DDPM) (Ho et al., 2020) to progressively guide noisy samples toward realistic dance motions, conditioned on the multi-modal embedding $E_{F}$ from the previous Multi-modal Fusion stage.
We model dance generation as a Markovian diffusion process, in which a sequence of noisy latent representations is gradually denoised to reveal the desired motion. The forward diffusion process begins with a clean motion sequence $\mathcal{X}$ and iteratively adds Gaussian noise over $T$ timesteps, eventually transforming it into a nearly random noise sample $z_{T}$. Formally, at each timestep $t$, the forward process is defined as:
(3) $\mathit{q(z_{t}|z_{t-1})}=\mathcal{N}(\sqrt{\mathit{1-\beta_{t}}}\mathit{z_{t-1}},\beta_{t}\mathit{I}),$
where $\beta_{t}\in(0,1)$ is a variance schedule parameter that controls the amount of noise added at each step. By the final timestep $T$, the noisy sequence $z_{T}$ closely approximates a standard Gaussian distribution $\mathcal{N}(0,\mathit{I})$.
To recover the original motion from this noisy representation, we employ a reverse diffusion process. This reverse process begins with the noisy sample $z_{T}$ and iteratively denoises it, guided by the conditional embedding $E_{F}$, to approximate the original sequence $\mathcal{X}$. At each step $t$, we predict the previous timestep $z_{t-1}$ from $z_{t}$ using:
(4) $\mathit{p_{\theta}(z_{t-1}|z_{t},E_{F})}=\mathcal{N}(\mathit{z_{t-1}};\mu_{\theta}(z_{t},t,E_{F}),\sigma_{t}^{2}\mathit{I}),$
where $\mu_{\theta}(z_{t},t,E_{F})$ represents the predicted mean for $z_{t-1}$ based on the model parameters $\theta$, and $\sigma_{t}$ is a learned variance that adjusts the noise schedule during the denoising process.
We adopt the simplified training objective from DDPM (Ho et al., 2020), minimizing the mean squared error (MSE) between the model’s predictions and the true motion data across all timesteps. This objective encourages the model to reconstruct realistic dance sequences from noisy inputs while aligning with the semantic and rhythmic cues encoded in $E_{F}$. Formally, the training objective is defined as:
(5) $\mathcal{L}_{\mathrm{simple}}=\mathbb{E}_{\mathbf{x},t}[\parallel\mathbf{x}-\mathbf{\hat{x}}(z_{t},t,E_{F})\parallel^{2}_{2}],$
where $\mathbf{\hat{x}}(z_{t},t,E_{F})$ denotes the reconstructed motion at timestep $t$ conditioned on the multi-modal embedding $E_{F}$.
3.6. Loss Functions

### v2

@@ Abstract
Music-to-dance generation aims to synthesize human dance motion conditioned on music input. Despite recent progress, significant challenges remain due to the semantic gap between music and dance motion, as music offers only abstract cues, but lacks explicit physical movement descriptions. The challenge is further amplified by the scarcity of paired music and dance data, which restricts the model’s ability to learn diverse dance patterns. These limitations highlight the need for additional semantic guidance beyond the musical signal. In this paper, we propose DanceChat, a novel framework that leverages a Large Language Model (LLM) as a choreographer to generate high-level textual instructions from structured music descriptions. These instructions serve as semantic guidance to bridge the gap between music and motion. DanceChat integrates music, beat, and text features into a unified representation, and employs a diffusion-based motion generator trained with a proposed multi-modal alignment loss. Extensive experiments on AIST++ dataset show that DanceChat outperforms state-of-the-art methods both qualitatively and quantitatively.
1 Introduction
@@
Dance has long served as a powerful medium of expression in cultures, playing a central role in rituals, social connections, and artistic activities (Carroll and Moore 2008). In the digital age, dance flourishes on social media, allowing people to share creativity and connect globally. Recent advances in deep learning have enabled computational exploration of this art form, particularly through music-to-dance generation, where models automatically create dance movements synchronized to music. This technology has impactful applications in entertainment, enabling realistic virtual idols and 3D characters in virtual reality. Beyond entertainment, automated dance generation is valuable in fields like robotics, digital human, and physical rehabilitation, where it helps create lifelike movement simulations and supports human-computer interaction.
Figure 1: Our approach draws inspiration from real-world dance learning, where learners rely on choreographers to interpret music. DanceChat leverages LLMs as pseudo-choreographers to translate music into textual instructions, bridging the semantic gap between music and motion. While the mapping from music to motion is inherently one-to-many and ill-posed, text offers a more structured and interpretable intermediary.
@@
Most existing music-to-dance generation approaches fall into two main categories: regression models (Huang et al. 2021; Li et al. 2021), which are typically designed to directly learn a mapping from input music to dance motions, and generative models (Siyao et al. 2022; Tseng, Castellon, and Liu 2023; Zhang et al. 2024a; Huang et al. 2024), which aim to learn a distribution over plausible dances conditioned solely on the input music. However, both approaches lack explicit movement instructions, and rely only on the abstract nature of musical elements. This creates a broad semantic gap between music and motion and complicates the task of generating physically plausible movements that align with the music’s genre and maintain aesthetic quality (Lee et al. 2019). Moreover, the scarcity of labeled dance data, due in part to the high cost of motion capture, makes it difficult for models to learn a well-generalized and diverse motion distribution. With limited training samples, the model struggles to capture the full variability of human dance, resulting in outputs that may lack stylistic diversity or fail to generalize to unseen music.
@@
These challenges highlight the need for advanced methods capable of bridging complex semantic gaps and injecting semantic-rich textual instructions without relying on additional annotations. Recent advancements in large language models (LLMs) have demonstrated remarkable success in capturing semantic relationships across various domains (Wu et al. 2023; Han et al. 2023; Doh et al. 2024; Gardner et al. 2024; Ding et al. 2024). They are trained on diverse content, including detailed discussions on music and dance, enabling them to interpret musical elements such as tempo, key, and chord progressions, as well as recognize movement styles, choreographic vocabulary, and dance semantics. This allows them to generate contextually appropriate descriptions that capture not only the structure and rhythm of a musical piece, but also align with plausible dance patterns. This inspires us to explore the potential of leveraging LLMs to enhance the diversity of music-to-dance generation.
@@
Building on these insights, our method adapts the capabilities of LLMs to address the semantic gap in music-driven dance generation. We propose a novel approach, named DanceChat, which conceptualizes the generation model as a dancing learner and leverages LLMs as a pseudo-choreographer (Fig. 1). Specifically, we extract structured musical descriptors, including key, tempo, and chord progressions alongside the dance genre context, and feed this information into LLM, which then generates corresponding textual dance instructions, serving as a pseudo-guide to condition the motion generation model during training. This helps reduce the semantic gap by offering interpretative cues that map abstract musical elements to concrete motion patterns.
@@
To further enhance cross-modal alignment, we explicitly extract beat and downbeat information from the input music as an additional rhythmic modality, which is known to correlate strongly with movement timing in dance. The original music, rhythmic cues, and LLM-generated textual instructions are then integrated through a multi-modal fusion module, enabling the model to form a richer and more coherent understanding of the music. By combining music, beat signals, and semantic instructions, DanceChat produces dance motions that are rhythmically aligned, physically plausible, and stylistically expressive.
@@ 2 Related Works
Figure 2: Overview of our approach. DanceChat consists of three main components: (a) Dance Instruction Generation module extracts music caption (tempo, key, chord, etc.) from the given music condition, (b) Multi-modal Fusion module integrates the encoded representations of music, beats, and text into a unified conditional embedding, and (c) Diffusion-based Motion Synthesis module uses this unified embedding to guide the generation of realistic dance movement.
2.1 3D Human Motion Generation
3D human motion generation has become a central problem in computer vision and graphics, with applications in animation, robotics, and virtual agents. Recent research has explored a wide range of conditioning modalities, including natural language, audio, scene layout, and control signals, aiming to improve realism, diversity, and controllability. Among these, text-conditioned motion generation has emerged as a promising direction, leveraging the semantic clarity and expressiveness of natural language to provide interpretable and flexible guidance. Diffusion-based approaches (Kim, Kim, and Choi 2023; Zhang et al. 2024b; Liang et al. 2024; Zhou et al. 2024) have demonstrated that structured textual inputs can guide the synthesis of diverse and coherent motion. This highlights the potential of language as a strong semantic prior for motion planning. Building on this idea, our work proposes to use language as a semantic bridge between music and motion. By translating musical attributes into textual instructions, we enable explicit, high-level guidance for music-to-dance generation, addressing the limitations of relying solely on music features.
2.2 Music-to-Dance Generation
@@
The Music-to-Dance generation task poses unique challenges due to the need for models to understand and generate dance sequences that are both rhythmically and stylistically aligned with music. Previous methods (Chan et al. 2019; Zhuang et al. 2022; Li et al. 2023; Tseng, Castellon, and Liu 2023; Yang, Yang, and Wang 2023; Tseng, Castellon, and Liu 2023; Huang et al. 2024) incorporate advanced audio-processing techniques with motion generation algorithms to enhance the flexibility and expressiveness of generated dances. Despite these advancements, existing approaches still face limitations in bridging the substantial semantic gap between music and dance, making the synchronization of dance sequences a significant challenge. Unlike text-to-motion generation where textual cues explicitly describe body movements, music provides only implicit cues through rhythm, tempo, and chord, which do not directly correspond to physical actions. This gap complicates the creation of dance sequences that are both well-aligned with the music and stylistically diverse.
2.3 Large Language Models
@@
Recent instruction-tuned LLMs (OpenAI 2023; Anthropic 2023; Touvron et al. 2023; DeepSeek 2024) have demonstrated strong capabilities in capturing contextual semantics, following user intent, and generalizing across domains. Inspired by these advancements, researchers have begun exploring LLMs for non-linguistic tasks, including motion generation (Zhou, Wan, and Wang 2024; Ribeiro-Gomes et al. 2024), where their ability to model temporal coherence and semantic structure enables fine-grained motion control. To effectively integrate LLMs into non-linguistic domains, prompt engineering has emerged as a key technique (Liu et al. 2021; Liu and Chilton 2022; Oppenlaender 2023). This involves crafting structured prompts that effectively communicate with models, directing the generation toward desired outcomes, and allowing users to control both the content and style of the outputs. In motion generation, prompt engineering enables models to translate high-level textual instructions into semantically coherent motion sequences, enhancing controllability and expressiveness (Wang et al. 2024). Building upon these developments, our work introduces a novel approach that leverages LLM-generated textual choreography instructions to bridge the semantic gap between music and dance.
3 Our Approach
3.1 Problem Definition
@@
The task of music-to-dance generation involves producing a dance motion sequence $\mathcal{X}=\{\mathit{x}^{1},\mathit{x}^{2},\dots,\mathit{x}^{N}\}$ from a given music track $\mathcal{M}$ with a duration of $\mathit{N}$ frames. Here, $\mathit{x}^{i}\in\mathbb{R}^{D}$ represents a human pose at the i-th frame, which is denoted as a D-dimensional vector. We present the dance sequences with 24-joint SMPL format (Loper et al. 2023), which represents each joint with 6-DOF rotation representation (Zhou et al. 2019) and a single root translation. Along with the binary contact labels $b\in\{0,1\}^{2\times 2=4}$ for the heel and the toe of each foot, the dimension $D=24\times 6+3+4=151$.
3.2 Overview
@@
Our model architecture is illustrated in Fig. 2 and consists of three main components: Dance Instruction Generation, Multi-modal Feature Extraction and Fusion, and Diffusion-based Motion Synthesis. The Dance Instruction Generation module (Fig. 2a) extracts music caption (tempo, beat, key, chord) from the given original music $\mathcal{M}$ and utilizes an LLM as a pseudo-choreographer to transform this information into dance motion instructions $\mathcal{T}$, providing textual guidance for motion generation. The Multi-modal Fusion module (Fig. 2b) then integrates the encoded representations of music $\mathcal{M}$, beats $\mathcal{B}$, and text $\mathcal{T}$ into a unified conditional embedding. This fused representation aligns music, beat, and movement instructions, creating a cohesive context for dance generation. Finally, the Diffusion-based Motion Synthesis module (Fig. 2c) uses this unified embedding to guide the generation of dance movements that capture the style and dynamics of the music.
3.3 Dance Instruction Generation
@@
This module is responsible for transforming music into high-level textual dance instructions, which serve as semantic guidance for the dance generation process. This module comprises two key steps: Music Caption Construction and LLM-based Instruction Generation.
Music Caption Construction. We adopt a structured captioning procedure to extract and represent key musical attributes from each input music clip. Building upon the framework proposed by Mustango (Melechovsky et al. 2023), we extract four common symbolic descriptors: tempo, chords, keys, and beat structure using a set of specialized tools. To further enhance clarity and generation consistency, we follow a controlled template-based sentence construction process, converting each musical attribute into interpretable textual music captions (e.g., ‘The chord sequence is G. The time signature is 4/4. The song is played at the pace of Moderato. The key of this song is G minor.’). These captions are appended to the input prompts to provide rich meta-level guidance for downstream instruction generation.
LLM-based Instruction Generation. To convert music into interpretable guidance for motion, we generate a textual dance instruction $\mathcal{T}$ from the music caption using an instruction-following LLM. The LLM acts as a pseudo choreographer, producing stylistically and rhythmically aligned instructions for each music clip. These instructions provide explicit descriptions of body movements, offering fine-grained guidance to the motion generator. To ensure consistency and controllability, we adopt a carefully designed prompt engineering strategy inspired by the format and motion description style of the HumanML3D dataset (Guo et al. 2022). As shown in Fig 3, we incorporate key elements such as clip duration and dance style. We also avoid directional words and overly fine-grained spatial cues to improve clarity and reduce ambiguity.
Figure 3: Details of our LLM-based textual instruction generation process. The prompt guides an LLM to generate HumanML3D-style dance instructions.
@@
Instruction Results. The output of this process is a set of natural language dance instructions that provide semantic cues for the dance generation model. For instance (Fig. 3), given a music track with a moderate tempo and a Hip-Hop style, the generated instruction might read: This person bounces with the knees while alternating shoulder pops and chest pops, adding a quick head nod on every beat. These instructions capture both the rhythmic alignment and stylistic characteristics of the music, bridging the semantic gap between music and motion.
3.4 Multi-modal Feature Extraction and Fusion